# Cost-Efficient Serving of LLM Agents via Test-Time Plan Caching

### v1

@@ Abstract
LLM-based agentic applications have shown increasingly remarkable capabilities in complex workflows but incur substantial costs due to extensive planning and reasoning requirements. Existing LLM caching techniques (like context caching and semantic caching), primarily designed for serving chatbots, are insufficient for agentic applications where outputs depend on external data or environmental contexts. We propose agentic plan caching, a novel approach that extracts, stores, adapts, and reuses structured plan templates from planning stages of agentic applications across semantically similar tasks to reduce the cost of serving. Unlike traditional semantic caching, our system extracts plan templates from completed agent executions at test-time, employs keyword extraction to match new requests against cached plans, and utilizes lightweight models to adapt these templates to task-specific plans with contexts. Evaluation across multiple real-world agentic applications shows that our system can reduce costs by 46.62% on average while maintaining performance, offering a more efficient solution for serving LLM-based agents that complements existing LLM serving infrastructures.
@@ 1 Introduction
Agentic applications based on Large Language Models (LLMs) have shown early promise in replicating human performance on a broad range of workflows, from coding [24, 26, 52] to web navigation [17, 21, 60] to open-ended research [3, 8]. Many of these LLM-based agents follow a two-stage pipeline that alternates between [54, 39]: (1) Plan – reasoning about what to do next, and (2) Act – executing those plans. While effective, these agents incur significant costs due to the complexity of workflows executed [29, 57] and need to interact with external tools and environments [38]. Specifically, the Plan stage is often implemented via test-time compute techniques  [11, 41] like chain-of-thought reasoning [46], which can require numerous LLM queries and access to expensive LLMs (e.g., reasoning models). This results in substantial costs for executing agentic workflows via APIs [13, 35] or locally [33].
To reduce LLM costs, methods have been developed to optimize responses to individual queries [30, 59]. In particular, caching has emerged as a popular approach, with two primary implementations: Context caching (e.g., KV cache reuse and prompt caching [18, 53, 51]) stores internal model states to speed up subsequent generations, while semantic caching [10, 40, 2] stores and reuses (input, output) pairs to accelerate the serving of queries that are similar to historical queries.
@@
To overcome these limitations, we propose agentic plan caching, which reduces the serving costs of LLM-based agents that follow the Plan-Act paradigm by adapting and reusing prior execution plans across semantically similar workflows. Our key insight is that the Plan stage, which incurs the majority of LLM compute cost, is often repeated (within or across workflows) despite yielding outputs that could be reused in future requests while maintaining performance. When an agent completes an execution of a workflow, we extract structured plan templates from the agent execution log. When a similar request arrives, we employ keyword extraction to identify the most important semantic target of the query, then match it against the cache to retrieve the most relevant plan template. Our approach differs from semantic caching by avoiding query-based cache lookups, which we found sub-optimal for agent applications. Whenever additional planning is required, we utilize a lightweight model to adapt the cached structured plan template into more detailed plans with task-specific contexts (e.g., fiscal year and company name in financial data-intensive reasoning [35]), rather than employing an expensive model.
Although several memory architectures have been proposed to help agents store and learn from past experiences  [42, 50, 44, 36], these efforts primarily focus on using such memories to improve the agent’s accuracy on completing workflows (e.g., with fewer hallucinations [9] or with higher task success rate [44]) rather than to reduce the cost of serving the agent. To our knowledge, the use of historical experiences to more efficiently serve LLM-based agents remains underexplored, particularly for applications where outputs depend on input data or environmental conditions external to the query itself.
We evaluate agentic plan caching on two applications and find that it reduces costs by 46.62% (on average) while maintaining 96.67% of optimal accuracy. The agentic plan caching we propose is compatible with existing LLM serving and agent frameworks, and can be used jointly with existing caching techniques as well.
@@
  3. 3.
    Evaluation of Caching Techniques: We evaluate our agentic plan caching system on top of real-world agent architecture and two datasets/benchmarks, and find that our approach can reduce cost by 46.62% (on average), while maintaining 96.67% of optimal application performance.
We open-source our code and artifact upon publication to facilitate further research in caching for efficient agentic AI applications.
@@ 2.1 Plan-Act Agents
The rise of large language models (LLMs) has driven the rapid expansion of agentic AI applications. Unlike single-model tasks like chatbots [15], math [22], or coding [14], these applications coordinate multiple models and queries to solve complex tasks, like data-intensive reasoning [35], software engineering [55, 45], web navigation [60], etc.
Many such agentic AI applications, like multi-agent systems [43, 20] and cloud-edge LLM systems [56, 35] , follow a two-stage pipeline loop (similar to the ReAct-agent loop [54]), as shown in Figure 1(a): (1) Plan and (2) Act. In the Plan stage, a planner LLM generates a strategy (e.g., task decomposition, information retrieval) that guides subsequent actions of acting LLMs. In the Act stage, the actor LLM acts accordingly based on devised plans and external context or environment, and passes down the response to the planner LLM for the next step.
However, due to the use of multiple LLMs and queries, especially with advanced models like reasoning or multimodal LLMs, these agentic applications can incur significant costs [28, 37], particularly in terms of token ingestion/generation. Optimizing these costs is crucial for scaling agentic AI applications.
@@ 2.2 LLM Caching: Methods and Limitations
Caching is one of the most widely-adopted techniques for reducing the serving cost of LLM applications. The goal of caching is to eliminate redundant computation. Context caching [18, 53, 51], also known as KV cache reuse or prompt caching, involves storing and reusing the key-value pairs generated during the prefill phase of LLM inference. Semantic caching [10, 40, 2], on the other hand, stores input-output pairs of previous LLM invocations. This relies on the fact that many prompts share similar underlying intents and thus expected outputs despite having different wording [40].
@@
Context caching relies on KV cache as the medium for storing and reusing knowledge [32, 18, 51]. These KV caches are inherently model-dependent and not easily transferable across different models [49, 31], since even identical text prompts produce model-specific KV caches. While this limitation is negligible for chatbots that consistently use a single model with the same system prompt, it becomes a problem for agentic AI applications that typically employ multiple LLMs across various processing stages.
@@
Semantic caching stores input-output pairs from previous LLM calls, assuming outputs depend solely on input prompts [10, 40]. While this holds for chatbots, many agentic AI applications are data-dependent: Outputs depend not only on input queries but also on external data (e.g., data-intensive reasoning [35]) or dynamic environments (e.g., web or GUI agents [60, 48, 47]). This dependency complicates the reuse of cached responses even when input prompts are semantically similar.
@@
Both context and semantic caching lack flexibility for handling slight variations in input. Context caching requires exact text matches. Semantic caching, while more accommodating, does not capture the transformation process from prompt to response. This could hinder adaptation to similar queries with minor differences (e.g., numeric values or variable names in mathematical reasoning [16], coding tasks [24]), a common challenge in agentic AI.
@@ 2.3 Related Work
Prior work has explored augmenting LLM agents with external memory to (1) reduce hallucinations through context-aware responses [9] and (2) enable complex, long-horizon tasks [44]. Some studies focus on defining memory formats [54] and managing memory efficiently [36, 50]. While our caching system can be adapted as a form of agent memory, it diverges by targeting serving cost reduction rather than enhanced capability, a largely unexplored area.
@@
Existing LLM serving engines like vLLM [30] and SGLang [59] optimize general query inference at scale through techniques such as KV cache management and request scheduling. Our approach is compatible with these systems, extending their capabilities to incorporate cost-effective caching for agentic AI scenarios.
3 The Agentic Plan Caching Framework
@@ 3.1 Overview of System Design
For a cache hit (Figure 2(a)), a small planner LM ("Small LM") adapts the retrieved plan template for the current execution by incorporating context-specific information (e.g., user information, environment variables). For a cache miss (Figure 2(b)), a large planner LM ("Large LM") generates a new plan from the input task query. The adapted or generated plan, along with the task context (e.g., external data or web/GUI environment), is then passed to the "Actor LM", which produces a response. The response is evaluated by the "Planner LM" to determine if further iterations are needed. If the task is complete, the final output is generated, concluding the agent’s execution.
Figure 2: Agentic Plan Caching Framework. We show: (a) cache hit workflow, (b) cache miss workflow, and (c) plan template generation upon cache miss.
@@ 3.2 Design Choices
A common method for identifying similar queries in a cache is to assess semantic or textual similarity, as seen in frameworks like GPTCache [10] which use embeddings for similarity searches. However, we find that query-based similarity matching, despite its popularity, is insufficient for detecting cache hits/misses for agentic plan caching. This is because it might overemphasize context-specific details (e.g., names of individuals or companies) rather than the broader intent of queries, which makes it difficult to establish an effective similarity threshold [40]. This frequently leads to elevated rates of false positives (irrelevant cache hits) or false negatives (missed reuse opportunities). In contrast, extracting keywords that reflect the higher-level intent of queries provides a more reliable indicator of whether two queries would result in similar agentic plans, as illustrated in Figure 3.
@@
Our system uses exact matches (between keywords) to minimize false positives. While fuzzy search [25] (identifying cache hits based on similar but not identical keywords) could handle approximate key similarities and is feasible to integrate, we opted against it and leave it for future exploration for two main reasons: (1) Determining fuzzy matches based on semantic or textual similarity of keywords would reintroduce challenges faced by semantic caching, such as setting effective similarity thresholds, and (2) although lightweight LMs could potentially enable fuzzy matching, cache lookups must remain fast and cost-effective, particularly in low-hit-rate scenarios.
@@
One naive approach to reuse historical experience is to cache and reuse past agent execution logs (containing all inputs and outputs from planner and actor LMs) as in-context learning examples for the small planner LM. However, in our experiments (§4.2), we find that small planner LMs, usually based on small language models (e.g., we use LLaMa-3.2-8B), struggle to handle long-context and unfiltered agent execution logs even when containing reusable plan information. This motivates us to filter agent execution logs into high-quality plan templates, and re-adapt them so that small planner LMs can better take advantage of their information.
4 Results
We evaluate our agentic plan caching framework across two datasets focused on data-intensive reasoning, covering long-context financial data (FinanceBench [23]) as well as tabular mathematical data (Tabular Math Word Problems [34]).
@@ 4 Results
    Reduced Cost: Agentic plan caching reduces LLM serving costs by an average of 46.62% (§4.2).
@@
    High Accuracy: We maintain 96.67% of application-level performance compared to the accuracy-optimal baseline (§4.2).
@@ 4.1 Experiment Setup
Our agentic plan caching system is built on the Minion architecture (Figure 1(a)) from the Minions project [35], a sequential Plan-Act LLM framework that can be readily generalized. The Minion architecture is composed of a large (cloud-hosted) planner LM for reasoning and task decomposition and a smaller (locally hosted) actor LM with access to additional context for plan execution. Given a task query, the planner LM and the actor LM collaborate iteratively (as in Figure 1(a)) to produce a final output. We set the maximum number of iterations to be 10.
We evaluate our system on two data-intensive reasoning benchmarks: FinanceBench [23] (financial and long-context data reasoning) and Tabular Math Word Problems [34] (tabular and mathematical data reasoning). Implementation details and dataset specifics are provided in the Appendix.
@@
We assess application-level performance using GPT-4o as the evaluation model, as LLM-based evaluation is more effective than exact matches or F1 scores for numeric evaluation and long-form responses [12, 19, 58]. Cost is calculated based on input/output tokens and the latest API pricing from commercial LLM providers (OpenAI API [7] and TogetherAI API [1]). Additional details on evaluation models, prompts, and API pricing are included in the Appendix.
@@
For the main results (§4.2), we use GPT-4o [5] as the planner LM and LLaMa-3.2-8B [6] as both the small planner LM and actor LM. For keyword extraction and cache generation, we use GPT-4o-mini [4]. To demonstrate broader applicability, we include a sensitivity analysis with a wider range of models in the Appendix.
@@
    Semantic Caching: We implement a query-level semantic caching method based on previous work [10, 40]. Following the approach of GPTCache [10], [^1] We do not use the official GPTCache release as it (1) lacks support for post-GPT-4 OpenAI models and (2) relies on a deprecated version of the OpenAI API. we cache and reuse responses to individual queries, determining cache hits based on query-level similarity.
@@
    Full-History Caching (discussed in §3.2): Inspired by knowledge caching in retrieval-augmented generation [53, 27], this baseline caches the complete agent execution log, including inputs and outputs of all LLM agent components. Cache hits are determined by keyword-level similarity. Upon a hit, the cached execution log is used as an in-context example for the small planner LM to generate new plans.
4.2 Results Analysis
Main Results
Agentic plan caching reduces cost by 46.62% on average while maintaining 96.67% of application-level performance compared to the accuracy-optimal baseline. We note that:

### v2

# Agentic Plan Caching: Test-Time Memory for Fast and Cost-Efficient LLM Agents

@@ Abstract
LLM-based agent applications have shown increasingly remarkable capabilities in complex workflows but incur substantial costs and latency due to extensive planning and reasoning requirements. Existing LLM caching techniques (like context caching and semantic caching), primarily designed for serving chatbots, are insufficient for agent applications where outputs depend on external data and environmental contexts. We propose Agentic Plan Caching (APC), a novel test-time memory that extracts, stores, adapts, and reuses structured plan templates from planning stages of agent applications across semantically similar tasks to reduce the cost and latency of serving. Unlike traditional semantic caching, our system extracts plan templates from completed agent executions at test-time, employs keyword extraction to match new requests against cached plans, and utilizes lightweight models to adapt these templates to task-specific plans with contexts. Evaluation across multiple real-world agent applications shows that our system can reduce costs by 50.31% and latency by 27.28% on average while maintaining performance, offering a more efficient solution for serving LLM-based agents that complements existing LLM serving infrastructures.
@@ 1 Introduction
Agent applications based on Large Language Models (LLMs) have shown early promise in replicating human performance on a broad range of workflows, from coding [jain2024livecodebench, jimenez2023swe, yang2024swe] to web navigation [he2024webvoyager, zhou2023webarena] to open-ended research [gemini-deep-research, openai-deep-research] to social interactions [park2023generative, xie2024ai]. Many of these LLM-based agents follow a two-stage pipeline, often referred to as the ReAct-agent loop [yao2023react, sarukkai2025self], that alternates between: (1) Plan – reasoning about what to do next, and (2) Act – executing those plans. While effective, these agents incur significant costs due to the complexity of executed workflows [zhang2024chain, kwa2025measuring] and need to interact with external tools and environments [patil2024gorilla]. Specifically, the Plan stage is often implemented via test-time compute techniques  [brown2024large, snell2024scaling] like chain-of-thought reasoning [wei2022chain], which can require numerous LLM queries and access to expensive LLMs (e.g., reasoning models). This results in substantial costs for executing agentic workflows via APIs [chen2023frugalgpt, narayan2025minions] or locally [liu2024mobilellm].
To reduce LLM costs, methods have been developed to optimize responses to individual queries [kwon2023efficient, zheng2024sglang]. In particular, caching has emerged as a popular approach, with two primary implementations: Context caching (e.g., KV cache reuse and prompt caching [gim2024prompt, yang2025kvlink, yao2025cacheblend]) stores internal model states to speed up subsequent generations, while semantic caching [bang2023gptcache, schroeder2025adaptive, amazon-semantic-cache] stores and reuses (input, output) pairs to accelerate the serving of queries that are similar to historical queries.
@@
To overcome these limitations, we propose Agentic Plan Caching (APC), a novel test-time memory that reduces the serving costs of LLM-based agents that follow the Plan-Act paradigm by adapting and reusing prior execution plans across semantically similar workflows. Our key insight is that the Plan stage, which incurs the majority of LLM compute cost, is often repeated (within or across workflows) despite yielding outputs that could be reused in future requests while maintaining performance. When an agent completes an execution of a workflow, we extract structured plan templates from the agent execution log. When a similar request arrives, we employ keyword extraction to identify the most important semantic target of the query, then match it against the cache to retrieve the most relevant plan template. Our approach differs from semantic caching by avoiding query-based cache lookups, which we found sub-optimal for agent applications. Whenever additional planning is required, we utilize a lightweight model to adapt the cached structured plan template into more detailed plans with task-specific contexts (e.g., fiscal year and company name in financial data-intensive reasoning [narayan2025minions]), rather than employing an expensive model.
Although several memory architectures have been proposed to help agents store and learn from past experiences  [sumers2023cognitive, packer2023memgpt, wang2024agent, xu2025mem, zhang2025agentic], these efforts primarily focus on using such memories to improve the agent’s accuracy on completing workflows (e.g., with fewer hallucinations [amazon-agent-cache] or with higher task success rate [wang2024agent]) rather than to reduce the cost of serving the agent. To our knowledge, the use of historical experiences to more efficiently serve LLM-based agents remains underexplored, particularly for applications where outputs depend on input data or environmental conditions external to the query itself.
We evaluate agentic plan caching on five diverse agent workloads and find that it reduces costs by 50.31% and latency by 27.28% (on average) while maintaining 96.61% of optimal accuracy. The agentic plan caching we propose is compatible with existing LLM serving and agent frameworks, and can be used jointly with existing caching techniques as well.
@@
  • 3.
    Evaluation of Caching Techniques: We evaluate our agentic plan caching system on top of real-world agent architecture and five datasets/benchmarks, and find that our approach can reduce cost by 50.31% and latency by 27.28% (on average), while maintaining 96.61% of optimal application performance.
@@ 2.1 Plan-Act Agents
The rise of large language models (LLMs) has driven the rapid expansion of agentic AI applications. Unlike single-model tasks like chatbots [chiang2024chatbot], math [hendrycks2021measuring], or coding [chen2021evaluating], these applications coordinate multiple models and queries to solve complex tasks, like data-intensive reasoning [narayan2025minions], software engineering [zhang2025adaptive, wei2024improving], web navigation [zhou2023webarena], etc.
Many such agentic AI applications, like multi-agent systems [wang2024mixture, guo2024large] and cloud-edge LLM systems [zhang2024caravan, narayan2025minions] , follow a two-stage pipeline loop (similar to the ReAct-agent loop [yao2023react]), as shown in Figure 1(a): (1) Plan and (2) Act. In the Plan stage, a planner LLM generates a strategy (e.g., task decomposition, information retrieval) that guides subsequent actions of acting LLMs. In the Act stage, the actor LLM acts accordingly based on devised plans and external context or environment, and passes down the response to the planner LLM for the next step.
However, due to the use of multiple LLMs and queries, especially with advanced models like reasoning or multimodal LLMs, these agentic applications can incur significant costs [jin2024efficient, pan2025specreason], particularly in terms of token ingestion/generation. Optimizing these costs is crucial for scaling agentic AI applications.
@@ 2.2 LLM Caching: Methods and Limitations
Caching is one of the most widely-adopted techniques for reducing the serving cost of LLM applications. The goal of caching is to eliminate redundant computation. Context caching [gim2024prompt, yao2025cacheblend, yang2025kvlink, xie2025strata], also known as KV cache reuse or prompt caching, involves storing and reusing the key-value pairs generated during the prefill phase of LLM inference. Semantic caching [bang2023gptcache, schroeder2025adaptive, amazon-semantic-cache], on the other hand, stores input-output pairs of previous LLM invocations. This relies on the fact that many prompts share similar underlying intents and thus expected outputs despite having different wording [schroeder2025adaptive].
@@
Context caching relies on KV cache as the medium for storing and reusing knowledge [liu2024cachegen, gim2024prompt, yang2025kvlink]. These KV caches are inherently model-dependent and not easily transferable across different models [wu2025know, liu2024droidspeak], since even identical text prompts produce model-specific KV caches. While this limitation is negligible for chatbots that consistently use a single model with the same system prompt, it becomes a problem for agentic AI applications that typically employ multiple LLMs across various processing stages.
@@
Semantic caching stores input-output pairs from previous LLM calls, assuming outputs depend solely on input prompts [bang2023gptcache, schroeder2025adaptive]. While this holds for chatbots, many agentic AI applications are data-dependent: Outputs depend not only on input queries but also on external data (e.g., data-intensive reasoning [narayan2025minions]) or dynamic environments (e.g., web or GUI agents [zhou2023webarena, wornow2024wonderbread, wornow2024automating]). This dependency complicates the reuse of cached responses even when input prompts are semantically similar.
@@
Both context and semantic caching lack flexibility for handling slight variations in input. Context caching requires exact text matches. Semantic caching, while more accommodating, does not capture the transformation process from prompt to response. This could hinder adaptation to similar queries with minor differences (e.g., numeric values or variable names in mathematical reasoning [cobbe2021training], coding tasks [jain2024livecodebench]), a common challenge in agentic AI.
@@ 2.3 Related Work
Prior work has explored augmenting LLM agents with external memory to (1) reduce hallucinations through context-aware responses [amazon-agent-cache] and (2) enable complex, long-horizon tasks [wang2024agent]. Some studies focus on defining memory formats [yao2023react] and managing memory efficiently [packer2023memgpt, xu2025mem]. While our caching system can be adapted as a form of agent memory, it diverges by targeting serving cost reduction rather than enhanced capability, a largely unexplored area.
@@
Existing LLM serving engines like vLLM [kwon2023efficient] and SGLang [zheng2024sglang] optimize general query inference at scale through techniques such as KV cache management and request scheduling. Our approach is compatible with these systems, extending their capabilities to incorporate cost-effective caching for agentic AI scenarios.
Case-Based Planning
Case-Based Planning (CBP) [bergmann1996role, spalzzi2001survey, borrajo2015progress] is a problem-solving paradigm where new plans are produced by adapting previously solved cases rather than constructing plans from scratch. By exploiting similarities between past and current situations, CBP supports efficient plan reuse, continual learning, and incremental refinement. While our work shares the high-level intuition of "plan storage and reuse", it targets a fundamentally different setting from classic symbolic CBP: LLM-based, neural agents performing open-ended natural-language actions. Specifically, our system (1) automatically extracts reusable plan templates at test time from unconstrained LLM generations, instead of relying on hand-crafted symbolic plans, and (2) leverages these stored plans for cost-efficient inference in LLM agents.
3 The Agentic Plan Caching (APC) Framework
@@ 3.1 Overview of System Design
Figure 2: Agentic Plan Caching Framework. We show: (a) cache hit workflow, (b) cache miss workflow, and (c) plan template generation for new cache entries.
@@
For a cache hit – Figure 2(a) – , a small planner LM ("Small LM") adapts the retrieved plan template for the current execution by incorporating context-specific information (e.g., user information, environment variables). For a cache miss – Figure 2(b) – , a large planner LM ("Large LM") generates a new plan from the input task query. The adapted or generated plan, along with the task context (e.g., external data or web/GUI environment), is then passed to the "Actor LM", which produces a response. The response is evaluated by the "Planner LM" to determine if further iterations are needed. If the task is complete, the final output is generated, concluding the agent’s execution.
@@ 3.2 Design Choices
A common method for identifying similar queries in a cache is to assess semantic or textual similarity, as seen in frameworks like GPTCache [bang2023gptcache] which use embeddings for similarity searches. However, we find that query-based similarity matching, despite its popularity, is insufficient for detecting cache hits/misses for agentic plan caching. This is because it might overemphasize context-specific details (e.g., names of individuals or companies) rather than the broader intent of queries, which makes it difficult to establish an effective similarity threshold [schroeder2025adaptive]. This often results in a high number of false positives (irrelevant cache hits) or false negatives (missed reuse opportunities). In contrast, extracting keywords that reflect the higher-level intent of queries provides a more reliable indicator of whether two queries would result in similar agentic plans, as illustrated in Figure 3.
@@
Our system uses exact matches (between keywords) to minimize false positives. While fuzzy search [ji2009efficient] (identifying cache hits based on similar but not identical keywords) could handle approximate key similarities and is feasible to integrate, we opted against it and leave it for future exploration for two main reasons: (1) Determining fuzzy matches based on semantic or textual similarity of keywords would reintroduce challenges faced by semantic caching, such as setting effective similarity thresholds, and (2) although lightweight LMs could potentially enable fuzzy matching, cache lookups must remain fast and cost-effective, particularly in low-hit-rate scenarios. We present an in-depth analysis of the overhead and scalability of fuzzy matching in §4.4.
@@
One naive approach to reuse historical experience is to cache and reuse past agent execution logs (containing all inputs and outputs from planner and actor LMs) as in-context learning examples for the small planner LM. However, in our experiments (§4.2), we find that small planner LMs, usually based on small language models (e.g., we use LLaMa-3.1-8B), struggle to handle long-context and unfiltered agent execution logs even when containing reusable plan information. This motivates us to filter agent execution logs into high-quality plan templates, and re-adapt them so that small planner LMs can better take advantage of their information.
4 Results and Evaluation
We evaluate our agentic plan caching (APC) framework on five agent workloads that span a diverse range of data-intensive reasoning and agentic capabilities. These include long-context data reasoning (FinanceBench [islam2023financebench], QASPER [dasigi2021dataset]), mathematical reasoning (Tabular Math Word Problems [lu2022dynamic], AIME 2024 and 2025 [aime-dataset]), and multi-step agentic reasoning and tool use (GAIA [mialon2023gaia]).
@@ 4 Results and Evaluation
    Reduced Cost: APC reduces agent serving costs by an average of 50.31% (§4.2).
@@
    Reduced Latency: APC reduces agent serving latency by an average of 27.28% (§4.3).
  • $\bullet$
    High Accuracy: APC maintains 96.61% of application-level performance compared to the accuracy-optimal baseline (§4.2).
@@ 4.1 Experiment Setup
Our agentic plan caching system is built on the Minion architecture (Figure 1(a)) from the Minions project [narayan2025minions], a sequential Plan-Act LLM framework that can be readily generalized. The Minion architecture is composed of a large (cloud-hosted) planner LM for reasoning and task decomposition and a smaller (locally hosted) actor LM with access to additional context for plan execution. Given a task query, the planner LM and the actor LM collaborate iteratively (as in Figure 1(a)) to produce a final output. We set the maximum number of iterations to be 10.
While we adopt this architecture for clarity, APC is not restricted to Minion-like agent architecture; we demonstrate its integration into other agent architectures and report end-to-end results in §4.2. Additional implementation details and dataset specifications are provided in the Appendix.
@@
We assess application-level performance using GPT-4o as the evaluation model, as LLM-based evaluation is more effective than exact matches or F1 scores for numeric evaluation and long-form responses [chen2025xverify, goldie2025synthetic, zheng2023judging]. Cost is calculated based on input/output tokens and the latest API pricing from commercial LLM providers (OpenAI API [openai-api] and TogetherAI API [together-api]). Additional details on evaluation models, prompts, and API pricing are included in the Appendix.
@@
For the main results (§4.2), we use GPT-4o [gpt-4o] as the planner LM and LLaMa-3.1-8B [llama-3.2] as both the small planner LM and actor LM. For keyword extraction and cache generation, we use GPT-4o-mini [gpt-4o-mini]. To demonstrate broader applicability, we include a sensitivity analysis with a wider range of models in the Appendix.
@@
    Semantic Caching: We implement a query-level semantic caching method based on previous work [bang2023gptcache, schroeder2025adaptive]. Following the approach of GPTCache [bang2023gptcache] [^1] We do not use the official GPTCache release as it (1) lacks support for post-GPT-4 OpenAI models and (2) relies on a deprecated version of the OpenAI API., we cache and reuse responses to individual queries, determining cache hits based on query-level similarity. We set similarity thresholds to be 80%, 85%, and 90%; a lookup is considered a hit if the query-level similarity is above this threshold.
@@
    Full-History Caching (discussed in §3.2): Inspired by knowledge caching in retrieval-augmented generation [yao2025cacheblend, jin2024ragcache], this baseline caches the complete agent execution log, including inputs and outputs of all LLM agent components. Cache hits are determined by keyword-level similarity. Upon a hit, the cached execution log is used as an in-context example for the small planner LM to generate new plans.
4.2 Main Results
As shown in Figure 4 and Table 1, agentic plan caching reduces cost by 50.31% on average while maintaining 96.61% of application-level performance compared to the accuracy-optimal baseline. We note that: